# Fate of transient order parameter domain walls in ultrafast experiments

### v1

@@ Abstract
Stability of the interface— In addition to the motion within the interface, the defects also move in the $z$-direction due to the thermal noise $\eta$. From Eq. (1), the $z$-direction motion of a defect is a Brownian motion given by $d(z/\xi_{0})/d(\gamma t)=H_{z}$ with a random force $H_{z}$ dependent on the defect size. For a typical defect with size $L$, the random force for a vortex satisfies
@@
Here $h_{z}(L)=\int d^{2}\bm{r}\ (\partial_{z}\theta)^{2}$ where $\theta(\bm{r})$ is the profile of a vortex with size $L$ discussed below Eq. (7), see Sec. III B of Ref. [29]. For the $U(1)$ symmetric case as well as for the $L\ll\xi_{0}/\sqrt{2\beta}$ case with a nonzero $\beta$, one has $h_{z}(L)\sim\ln(L/\xi)$, while it is hard to evaluate $h_{z}(L)$ for the other cases. Eq. (9) implies that as long as $\zeta$ is small, the $z$-direction motion of the defects is negligible and the interface is stable.
Effects of the inertia term— So far, we have clarified the decaying process of the domain wall in terms of the “model A” dynamics, Eq. (1). The “model A” dynamics neglects the inertia of the order parameter, the $\partial_{t}^{2}\psi$ term, while the inertia is crucial for the inversion of the order parameter by the pump pulse that generates the domain wall [25, 26, 27, 28, 29]. Thus, we now proceed to add the inertia term $\gamma_{0}^{-2}\partial_{t}^{2}\psi$ to the left-hand side of Eq. (1), and discuss how the decaying processes of the domain wall are modified. Note that microscopic derivations of the relevant TDGL theory [40, 48] gives $\gamma,\,\gamma_{0}\sim T$ for pure electronic density-wave orders arising from Fermi surface nesting, see Ref. [40] and Appendix E of Ref. [48], while a mixing with lattice distortions may significantly increase the inertia.
In the first-stage dynamics, a small $|\alpha|$ suppresses the inertia term more effectively than the damping term. Thus, as long as the temperature is close to $T_{\text{c}}$ where $|\alpha|\ll 1$, the inertia term does not change the qualitative picture of the first-stage dynamics, modifying only the coefficients. Specifically, the time-dependent correlation length in Eqs. (4,6) is renormalized by a factor $1/\sqrt{\kappa}$ with $\kappa=\sqrt{1+4(|\alpha|-2\beta)\gamma^{2}/\gamma_{0}^{2}}$, see Sec. IV of Ref. [29].
@@
see Sec. IV of Ref. [29]. Note that all the results of $L(t)$ satisfying Eq. (5) obtained in this work satisfy Eq. (10) in the long-time limit, when the term $m(L)d^{2}L/dt^{2}$ is much smaller than the other two terms.
@@
Discussion—Our theory enriches the mechanism of topological defect generation in ultrafast experiments [50, 51, 52, 11, 24, 53]. It applies to a wide class of systems with order parameters breaking a (quasi) continuous symmetry, such as charge [54, 55, 45, 22, 56], spin [41], excitonic order [57, 58, 46], and superconductivity [39, 42, 40, 16].
@@
In some experiments for 3D systems, a strong pump fluence could generate multiple domain walls at different depths in the $z$ direction with a typical inter-wall distance $L_{z}$ [26, 25, 27]. In this paper, we focused on the decaying dynamics of an individual domain wall neglecting the interaction between the walls. This is appropriate as long as the inter-wall distance $L_{z}$ is much larger than the typical size $L$ of the topological defects on the walls, such that each wall evolves independently with a diffusive growth of in-plane correlation length $L$ while $L_{z}$ does not grow. In other words, the measured correlation-length dynamics depends on the direction of the measurement. This might be relevant to a recent pump-probe experiment that reports anomalous subdiffusive coarsening dynamics of the CDW in $\text{LaTe}_{3}$ [36].
@@
Similar physics also applies to systems with two competing orders [16, 29]. For example, a real field $\psi$ breaking a $Z_{2}$ symmetry (e.g., commensurate CDW) competes with a complex field $\Psi=\Psi_{1}+i\Psi_{2}$ breaking a $U(1)$ symmetry (e.g., superconductivity or incommensurate CDW) [55, 54] such that $\psi$ dominates and $\Psi$ is completely suppressed in the equilibrium. If a pump pulse generates a domain wall of $\psi$, the fluctuations of $\Psi$ may become unstable on the domain wall and grow into a random domain structure with increasing correlation length, see Sec. V of Ref. [29].
@@
Acknowledgment— We thank Andrew J. Millis, Mariano Trigo, Alfred Zong, Ke Liu, Yeyang Zhang, and Zhenyu Xiao for helpful discussions. Lingxian Kong and Ryuichi Shindou acknowledge the supports from the National Basic Research Programs of China (No. 2019YFA0308401) and the National Natural Science Foundation of China (No. 11674011 and No. 12074008). Zhiyuan Sun acknowledges the support of the startup grant from the State Key Laboratory of Low-Dimensional Quantum Physics at Tsinghua University.

### v2

@@ Results
Stability of the interface
In addition to the motion within the interface, the defects also move in the $z$-direction due to the thermal noise $\eta$. From Eq. (1), the $z$-direction motion of a defect is a Brownian motion given by $d(z/\xi_{0})/d(\gamma t)=H_{z}$ with a random force $H_{z}$ dependent on the defect size. For a typical defect with size $L$, the random force for a vortex satisfies
@@
Here $h_{z}(L)=\int d^{2}\bm{r}\ (\partial_{z}\theta)^{2}$ where $\theta(\bm{r})$ is the profile of a vortex with size $L$ discussed below Eq. (7), see the Supplementary Note 3. For the $U(1)$ symmetric case as well as for the $L\ll\xi_{0}/\sqrt{2\beta}$ case with a nonzero $\beta$, one has $h_{z}(L)\sim\ln(L/\xi)$, while it is hard to evaluate $h_{z}(L)$ for the other cases. Eq. (9) implies that as long as $\zeta$ is small, the $z$-direction motion of the defects is negligible and the interface is stable.
Effects of the inertia term
So far, we have clarified the decaying process of the domain wall in terms of the “model A” dynamics, Eq. (1). The “model A” dynamics neglects the inertia of the order parameter, the $\partial_{t}^{2}\psi$ term, while the inertia is crucial for the inversion of the order parameter by the pump pulse that generates the domain wall [27, 28, 29, 30]. Thus, we now proceed to add the inertia term $\gamma_{0}^{-2}\partial_{t}^{2}\psi$ to the left-hand side of Eq. (1), and discuss how the decaying processes of the domain wall are modified. Note that microscopic derivations of the relevant TDGL theory [42, 52] gives $\gamma,\,\gamma_{0}\sim T$ for pure electronic density-wave orders arising from Fermi surface nesting, see Ref. [42] and Appendix E of Ref. [52], while a mixing with lattice distortions and electron-phonon scattering may significantly modify these parameters.
In the first-stage dynamics, a small $|\alpha|$ suppresses the inertia term more effectively than the damping term. Thus, as long as the temperature is close to $T_{\text{c}}$ where $|\alpha|\ll 1$, the inertia term does not change the qualitative picture of the first-stage dynamics, modifying only the coefficients. Specifically, the time-dependent correlation length in Eqs. (4,6) is renormalized by a factor $1/\sqrt{\kappa}$ with $\kappa=\sqrt{1+4(|\alpha|-2\beta)\gamma^{2}/\gamma_{0}^{2}}$, see the Supplementary Note 4.
@@
see the Supplementary Note 4. Note that all the results of $L(t)$ satisfying Eq. (5) obtained in this work satisfy Eq. (10) in the long-time limit, when the term $m(L)d^{2}L/dt^{2}$ is much smaller than the other two terms.
Estimation of parameters and timescales.
Phenomenological parameters $\gamma,\xi_{0},\zeta$, and $\alpha$ determine the timescale of the dynamics. In relevant CDW materials such as RTe3 (R = La, Sm, Tb, Ce), 1T-TaS2, and 1T-TiSe2 [54, 55, 56], the relaxation time is roughly $1/\gamma\sim 1$ ps from ultrafast experiments [16, 11, 57, 29, 27, 20]. From the expression $\xi_{0}\sim v_{\rm F}/T_{\rm c}$, $\xi_{0}$ is estimated as several nanometers for LaTe3 [22], while a similar value was obtained for 1T-TaS2 from the domain-wall width in scanning tunneling microscopy images [20]. Experiments are typically performed in 3D materials, where $T_{\rm c}^{D-1}/E_{\rm F}^{D-1}\sim 10^{-4}$ given that $E_{\rm F}\sim 10^{4}-10^{5}$ K for metals and $T_{\rm c}\sim 10^{2}-10^{3}$ K. Therefore, the Ginzburg parameter is estimated by $\zeta\sim 10^{-4}/\sqrt{|\alpha|}$ for a certain experimental temperature. For example, in room-temperature ($T\approx 300$ K) experiments for SmTe3 with $T_{\rm c}=416$ K or LaTe3 with $T_{\rm c}=670$[29, 11] , one has $\sqrt{|\alpha|}\sim 1$ and $\zeta\sim 10^{-4}$. The first-stage dynamics finishes at a crossover time $t_{\rm c}\sim 10\gamma^{-1}\sim 10$ ps. The second-stage dynamics will continue until $L(t)$ reaches a certain length scale $L_{\rm s}$ when the domain wall starts to interact with other domain walls or the sample surface. $L_{\rm s}$ is the domain-wall depth or the inter-domain-wall distance if multiple domain walls are generated. From the literature [29, 28], an estimation gives $L_{\rm s}/\xi_{0}\sim 10$ in 3D systems. Therefore, the scaling $L(t)\sim\xi_{0}\sqrt{\gamma t}$ gives a timescale $\sim 10^{2}$ ps for the second-stage dynamics.
Experimental scheme.
The CDW order induces a lattice distortion $\bm{b}(\bm{r})\propto\psi(\bm{r},t)e^{i\bm{Q}\cdot\bm{r}}$, where $\bm{Q}$ is the CDW wavevector lying in the $x-y$ plane, see the Supplementary Note 7. One could measure the evolution of the CDW order from the lattice structure factor in ultrafast x-ray scattering [38, 58, 39]. For example, in a recent ultrafast x-ray scattering experiment for LaTe3 which exhibits an incommensurate CDW, the material was excited by a 1.55 eV, 35 fs optical pump pulse with the fluence of 8 mJ/cm2 and probed by a 50 fs, 10 keV hard x-ray pulse [38]. To investigate the dynamics of the $z=0$ interface parametrized by $\bm{R}$, one could extract a structure factor defined by $S(\bm{R},t)\equiv\langle b^{\dagger}(\bm{0},t)b(\bm{R},t)\rangle\propto\langle\psi^{\dagger}(\bm{0},t)\psi(\bm{R},t)\rangle e^{i\bm{Q}\cdot\bm{R}}$, whose Fourier transform $\tilde{S}(\bm{K},t)$ has a peak at $\bm{K}=\bm{Q}$ with the peak intensity corresponding to the amplitude $|\psi|^{2}$ and the inverse of the full width at half maximum (IFWHM) corresponding to the correlation length $L(t)$. In the long-time regime of the first-stage dynamics, Eqs. (4)(6) predict the peak intensity to show an exponential growth and the IFWHM to exhibit a diffusive growth. In the second-stage dynamics, the IFWHM is expected to show a diffusive growth for 3D materials while the peak intensity has no significant growth since the growth of fluctuation magnitude has saturated.
Discussion
@@
Our theory enriches the mechanism of topological defect generation in ultrafast experiments [59, 60, 61, 11, 25, 62]. It applies to a wide class of systems with order parameters breaking a (quasi) continuous symmetry, such as charge order [63, 64, 46, 23, 57], spin order [43], excitonic order [48, 49, 19], and superconductivity [41, 44, 42, 17].
@@
In some experiments for 3D systems, a strong pump fluence could generate multiple domain walls at different depths in the $z$ direction with a typical inter-wall distance $L_{z}$ [28, 27, 29]. In this paper, we focused on the decaying dynamics of an individual domain wall neglecting the interaction between the walls. This is appropriate as long as the inter-wall distance $L_{z}$ is much larger than the typical size $L$ of the topological defects on the walls, such that each wall evolves independently with a diffusive growth of in-plane correlation length $L$ while $L_{z}$ does not grow. In other words, the measured correlation-length dynamics depend on the direction of the measurement. This might be relevant to a recent pump-probe experiment that reports anomalous sub-diffusive coarsening dynamics of the CDW in $\text{LaTe}_{3}$ [38].
@@
Optically generated domain walls may also appear and lead to interesting dynamics in other situations that warrant further investigation [65, 66, 67], e.g., the domain wall between superconducting orders with different chirality [68]. Another example is the system with competing orders [17]: a real field $\psi$ breaking a $Z_{2}$ symmetry (e.g., commensurate CDW) competes with a complex field $\Psi=\Psi_{1}+i\Psi_{2}$ breaking a $U(1)$ symmetry (e.g., superconductivity or incommensurate CDW) [64, 63] such that $\psi$ dominates and $\Psi$ is completely suppressed in equilibrium. If a pump pulse generates a domain wall of $\psi$, the fluctuations of $\Psi$ may become unstable on the domain wall and grow into a random domain structure with increasing correlation length. A preliminary study of a model with a quartic competing term $|\Psi|^{2}\psi^{2}$ is provided in Supplementary Note 5, while it is meaningful to perform a systematical study with various competing terms between multi-component orders.
Methods
Both analytical and numerical results in this work are obtained by the “model A” dynamics, Eq. (1). The results of the first-stage dynamics are derived from the linearized dynamical equation, Eq. (3). The analytical results of the second-stage dynamics are obtained from the force-balance equation, Eq. (5), which is derived from Eq. (1). In the numerical simulations, the “model A” dynamical equation was rescaled by
$\alpha$ $=\bar{\alpha}b^{2},\ \ \beta=\bar{\beta}b^{2},\ \ E_{\text{c}}=\bar{E_{\text{c}}}b^{-4},\ \ \psi=\bar{\psi}b,$
$\ \ \eta$ $=\bar{\eta}b^{3},\ \ \xi_{0}=\bar{\xi_{0}}b,\ \ \gamma=\bar{\gamma}b^{-2},$ (11)
with $b=\sqrt{|\alpha|}$ and $\bar{\alpha}=-1$. We further define $\tau=|\alpha|\gamma t$, $\tilde{\bm{r}}=\bm{r}/\xi$ and ~$\tilde{\nabla}=\partial/\partial\tilde{\bm{r}}$ and update $\bar{\psi}$ after each time step according to the dynamical equation given in the following form:
$\bar{\psi}(\tilde{\bm{r}},\tau+d\tau)-\bar{\psi}(\tilde{\bm{r}},\tau)=\Big{(}\tilde{\nabla}^{2}\bar{\psi}(\tilde{\bm{r}},\tau)+2\bar{\psi}(\tilde{\bm{r}},\tau)-2\frac{\beta}{|\alpha|}i\bar{\psi_{2}}(\tilde{\bm{r}},\tau)-2|\bar{\psi}(\tilde{\bm{r}},\tau)|^{2}\bar{\psi}(\tilde{\bm{r}},\tau)\Big{)}d\tau+\bar{\eta}(\tilde{\bm{r}},\tau)d\tau.$ (12)
The rescaled noise correaltor reads $\langle\bar{\eta}_{i}(\tilde{\bm{r}},\tau)\bar{\eta}_{i^{\prime}}(\tilde{\bm{r}}^{\prime},\tau^{\prime})\rangle=2\zeta\delta_{ii^{\prime}}\delta(\tilde{\bm{r}}-\tilde{\bm{r}}^{\prime})\delta(\tau-\tau^{\prime})$. $\bar{\eta}(\tilde{\bm{r}},\tau)d\tau$ are generated by $\bar{\eta}_{i}(\tilde{\bm{r}},\tau)d\tau=\sqrt{\frac{2\zeta d\tau}{\Pi_{j}dr_{j}}}n_{i}(\tilde{r},\tau)$ with $n_{i}(\tilde{r},\tau)$ being generated from independent standard normal distribution for each discretized $\tilde{\bm{r}}$ and $\tau$. The values of $2\zeta$ and the discretization step sizes, $dr_{j=x,y,z}$ and $d\tau$, are provided in the caption of Fig. 3. The second-order derivative, ~$\tilde{\nabla}^{2}\bar{\psi}$, is computed using the central difference method. A free boundary condition is applied along the $z$-direction, while periodic boundary conditions are imposed along the $x$- and $y$-directions.
Data availability statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Code availability statement
The code for the numerical simulation of this study is available from the corresponding author upon reasonable request.
Acknowledgments
@@
We thank Andrew J. Millis, Mariano Trigo, Alfred Zong, Shaofeng Duan, Qingzheng Qiu, Ke Liu, Yeyang Zhang, and Zhenyu Xiao for helpful discussions. Lingxian Kong and Ryuichi Shindou acknowledge the support from the National Basic Research Programs of China (No. 2019YFA0308401) and the National Natural Science Foundation of China (No. 11674011 and No. 12074008). Zhiyuan Sun acknowledges the support from the National Natural Science Foundation of China (Grants No. 12421004 and No. 12374291) and the State Key Laboratory of Low-Dimensional Quantum Physics at Tsinghua University.
Author contributions
Z.S. conceived and supervised the project. L.K. performed the calculations. L.K., R.S., and Z.S. co-wrote the manuscript.
Competing Interests
The authors declare no competing interests.